OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. A. y − 6 = one half(x + 3) B. y = one halfx − 6 C. y + 3 = one half(x − 6) D. x − 2y = 12

OpenStudy (anonymous):

B

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

@Mertsj @Hero @ash2326

OpenStudy (anonymous):

I don't need help now this is just for a mod ty

Vocaloid (vocaloid):

let (x1,y1) be the point (6,-3) , let m = slope plug in x1, y1, and slope into the equation y-y1 = m(x-x1)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

yes?

OpenStudy (anonymous):

The first Guy Icedragon50 gave me a void answer He came in and a second later he typed B and left

OpenStudy (mertsj):

Is B the correct answer?

OpenStudy (anonymous):

No

OpenStudy (mertsj):

What is the correct answer?

OpenStudy (anonymous):

Never got it but after I submitted my test after no further help all my questions were correct except this one...

OpenStudy (anonymous):

The guy can't even give me an explanation!

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mertsj):

Did you use the point slope form of the equation?

OpenStudy (anonymous):

Yes if I could @Vocaloid gave it to me first

OpenStudy (mertsj):

The directions clearly say to choose the point slope form of the equation.

OpenStudy (mertsj):